### v1

# Norm of the Cesàro operator between some spaces of analytic functions

Abstract. In this paper, we determine the exact norm of the Cesàro operator $\mathcal{C}$ on the Korenblum space $H^{\infty}_{\alpha}$ for $0<\alpha\leq\frac{1}{2}$ and on the logarithmically weighted space $H^{\infty}_{\alpha,\log}$ for $0<\alpha<1$. Moreover, we compute its norm when acting from $H^{\infty}_{\alpha,\log}$ to $H^{\infty}_{\alpha}$. Finally, we establish lower and upper bounds for the norm of $\mathcal{C}$ on the $\alpha$-Bloch space $\mathcal{B}^{\alpha}$ for $\alpha>1$, and from the Hardy space $H^{\infty}$ to $\mathcal{B}^{\alpha}$ for $\alpha\geq 1$.
Keywords. Operator norms, Cesàro operators, Korenblum spaces, Bloch spaces.
footnotetext: Corresponding author. E-mail addresses: slye@zust.edu.cn (S. Ye).
1. Introduction
Let $(a)=\{a_{k}\}_{k=0}^{\infty}$ be a sequence of complex numbers. The classical Cesàro operator acting on such sequences is defined as
$\mathcal{C}((a)):=\left(\frac{1}{n+1}\sum_{k=0}^{n}a_{k}\right)_{n=0}^{\infty}.$
As early as the 1920s, Hardy [8] and Landau [13] established the boundedness of this operator on the $\ell^{p}$ spaces.
The Cesàro operator can also be interpreted as an operator acting on functions defined on $\mathbb{D}$, the unit open disc in the complex plane $\mathbb{C}$. Specifically, for any analytic function $f\in H(\mathbb{D})$, if its Taylor expansion is $f(z)=\sum_{k=0}^{\infty}a_{k}z^{k},\ z\in\mathbb{D}$, then the Cesàro operator $\mathcal{C}:H(\mathbb{D})\to H(\mathbb{D})$ is given by
$\mathcal{C}(f)(z):=\sum_{n=0}^{\infty}\left(\frac{1}{n+1}\sum_{k=0}^{n}a_{k}\right)z^{n}=\int_{0}^{1}\frac{f(tz)}{1-tz}dt.$
The boundedness and compactness properties of the Cesàro operator have been extensively studied in complex and functional analysis [6, 11, 15, 23]. Early investigations into the boundedness of Cesàro operators on Hardy spaces $H^{p}$ for $1<p<\infty$ relied on Hardy’s work on Fourier series [10] and M. Riesz’s theorem on conjugate functions [7, Theorem 4.1]. Using the theory of composition operator semigroups, Siskakis [21] gave an alternative proof of the above result and further studied the case $p=1$ in [20]. A different proof for the case $p=1$ was provided by Giang [14]. Later, Miao [15] showed that the Cesàro operator is bounded on $H^{p}$ for all $0<p<1$.
[…]
However, there are relatively few works on the exact norm computation of the classical Cesàro operator. The main known results in this direction are due to Siskakis. In [21], he established that $\|\mathcal{C}\|_{H_{p}}=p$ for $p\geq 2$, while for $1\leq p<2$, the norm satisfies $p\leq\|\mathcal{C}\|_{H_{p}}\leq 2$. In [19], he showed that $\|\mathcal{C}\|_{A_{p}}=p/2$ for $p\geq 4$, and $p/2\leq\|\mathcal{C}\|_{A_{p}}\leq 2$ for $1\leq p<4$. In [6], Danikas and Siskakis also obtained $\|\mathcal{C}\|_{H^{\infty}\to\mathrm{BMOA}}=1+\pi/\sqrt{2}$.
In this article, we study the norm of $\mathcal{C}$ acting between certain spaces of analytic functions. Our paper is organized as follows. In Sect. 2, we introduce some notation. In Sect. 3, we determine the exact value of the norm of $\mathcal{C}$ on the Korenblum space $H^{\infty}_{\alpha}$ for $0<\alpha\leq\frac{1}{2}$, which is $\frac{1}{\alpha}$. In Sect. 4, for $0<\alpha<1$, we calculate the exact value of the norm from the logarithmically weighted Korenblum space $H^{\infty}_{\alpha,\log}$ to the Korenblum space $H^{\infty}_{\alpha}$. In Sect. 5, we calculate the exact value of the norm on the logarithmically weighted Korenblum space $H^{\infty}_{\alpha,\log}$. In Sect. 6, we obtain both the lower and upper bounds of the norm on $\alpha$-Bloch space $\mathcal{B}^{\alpha}$. In Sect. 7, we offer both the lower and upper bounds of the norm of the Cesàro operator from the Hardy space $H^{\infty}$ to $\alpha$-Bloch spaces, show that $\mathcal{C}:H^{\infty}_{\alpha}\to\mathcal{B}^{\alpha}$ is not bounded when $0<\alpha<1$.
2. Notation Preliminaries
Let $\mathbb{D}$ denote the open unit disk of the complex plane $\mathbb{C}$, and let $H(\mathbb{D})$ denote the set of all analytic functions in $\mathbb{D}$.
Recall that for $0<p<\infty$, the Hardy space $H^{p}$ consists of all analytic functions $f\in H(\mathbb{D})$ satisfying
$\|f\|_{H^{p}}=\sup_{0\leq r<1}M_{p}(r,f)<\infty,$
where
$M_{p}(r,f)=\left(\frac{1}{2\pi}\int_{0}^{2\pi}|f(re^{it})|^{p}dt\right)^{\frac{1}{p}},\quad 0<p<\infty;$
$M_{\infty}(r,f)=\sup_{0\leq t<2\pi}|f(re^{it})|.$
We refer to [7] for the notation and results regarding Hardy spaces.
For $0<\alpha<1$, the Korenblum space $H^{\infty}_{\alpha}$ is the space of all functions $f\in H(\mathbb{D})$ such that
$\|f\|_{H^{\infty}_{\alpha}}=\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}|f(z)|<\infty.$
Next, we present the definition of the weighted Korenblum space, which we introduced in reference [12], for $0<\alpha<1$, the logarithmically weighted Korenblum spaces $H^{\infty}_{\alpha,\log}$ as the set of all $f\in H(\mathbb{D})$ such that
$\|f\|_{H^{\infty}_{\alpha,\log}}\overset{def}{=}\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-|z|^{2}}|f(z)|<\infty.$
It is easily verified that $H^{\infty}\subsetneqq H^{\infty}_{\alpha,\log}\subsetneqq H^{\infty}_{\alpha}$.
For $0<\alpha<\infty$, the $\alpha$-Bloch space $\mathcal{B}^{\alpha}$ consists of those functions $f\in H(\mathbb{D})$ with
$\|f\|_{\alpha*}=\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}|f^{\prime}(z)|<\infty.$
It is easy to check that $\|\ \|_{\mathcal{{}^{\alpha*}}}$ is a complete semi-norm on $\mathcal{B^{\alpha}}$, and $\mathcal{B^{\alpha}}$ can be made into a Banach space by introducing the norm
$\|f\|_{\mathcal{B\alpha}}=|f(0)|+\|f\|_{\alpha*}.$
We can see that $\mathcal{B}^{1}$ is the classical Bloch space $\mathcal{B}$. We mention [17, 24] as general references for the classical Bloch space and the $\alpha$-Bloch spaces.
For an analytic function $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ on the unit disk $\mathbb{D}$, the image $\mathcal{C}(f)$ is also analytic on $\mathbb{D}$ and admits several equivalent representations(See [21]). In particular, it can be expressed as:
$\mathcal{C}(f)(z)$ $=\sum_{n=0}^{\infty}\left(\frac{1}{n+1}\sum_{k=0}^{n}a_{k}\right)z^{n}$
$=\int_{0}^{1}\frac{f(tz)}{1-tz}\,dt$
$=\frac{1}{z}\int_{0}^{z}\frac{f(\xi)}{1-\xi}\,d\xi.$ (2.1)
By a change of variable in the integral representation, the Cesàro operator can be rewritten in terms of a family of weighted composition operators. Specifically, we have:
$\mathcal{C}(f)(z)=\int_{0}^{\infty}S_{t}f(z)\,dt,$ (2.2)
where
$S_{t}f(z)=w_{t}(z)f(\phi_{t}(z)),\quad w_{t}(z)=\frac{e^{-t}}{1-(1-e^{-t})z},\quad\phi_{t}(z)=\frac{e^{-t}z}{1-(1-e^{-t})z}.$
Differentiating under the integral sign yields the derivative of $\mathcal{C}(f)$:
$\mathcal{C}(f)^{\prime}(z)=\int_{0}^{\infty}\left[\frac{e^{-t}(1-e^{-t})}{(1-(1-e^{-t})z)^{2}}f(\phi_{t}(z))+\frac{e^{-2t}}{(1-(1-e^{-t})z)^{3}}f^{\prime}(\phi_{t}(z))\right]dt.$ (2.3)
3. Norm estimates of the Cesàro operator $\|\mathcal{C}\|_{H_{\alpha}^{\infty}\rightarrow H^{\infty}_{\alpha}}$
In this section, we establish norm estimates for the Cesàro operator acting on the Korenblum space $H^{\infty}_{\alpha}$.
Theorem 3.1.
For $0<\alpha\leq\frac{1}{2}$, the Cesàro operator $\mathcal{C}$ is bounded on Korenblum space $H_{\alpha}^{\infty}$, and its norm satisfies
$\|\mathcal{C}\|_{H_{\alpha}^{\infty}\rightarrow H^{\infty}_{\alpha}}=\frac{1}{{\alpha}}.$
Proof.
First, we consider the lower bound of $\|\mathcal{C}\|_{H_{\alpha}^{\infty}\to H_{\alpha}^{\infty}}$. Let $0<\alpha<1$ and $z\in\mathbb{D}$. Define
$f_{\alpha}(z)=\frac{1}{(1-z^{2})^{\alpha}}.$
On one hand, we have the estimate
$\|f_{\alpha}\|_{H_{\alpha}^{\infty}}=\sup_{z\in\mathbb{D}}\frac{(1-|z|^{2})^{\alpha}}{|1-z^{2}|^{\alpha}}\leq\sup_{z\in\mathbb{D}}\frac{(1-|z|^{2})^{\alpha}}{(1-|z|^{2})^{\alpha}}=1.$
On the other hand, for $r\in(0,1)$, it holds that
$\lim_{r\to 1^{-}}|f_{\alpha}(r)|(1-r^{2})^{\alpha}=1,$
and we obtain $\|f_{\alpha}\|_{H_{\alpha}^{\infty}}=1$.
Now,
$\|\mathcal{C}\|_{H_{\alpha}^{\infty}\to H_{\alpha}^{\infty}}$ $\geq\frac{\|\mathcal{C}(f_{\alpha})\|_{H_{\alpha}^{\infty}}}{\|f_{\alpha}\|_{H_{\alpha}^{\infty}}}$
$=\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\left|\int_{0}^{\infty}S_{t}f_{\alpha}(z)dt\right|$
$=\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\left|\int_{0}^{\infty}\frac{e^{-t}}{1-(1-e^{-t})z}\cdot\frac{1}{(1-(\phi_{t}(z))^{2})^{\alpha}}dt\right|$
$\geq\sup_{0\leq r<1}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}}{\left(1-(1-2e^{-t})r\right)^{\alpha}}dt.$
Letting $r\to 1^{-}$, we obtain
$\lim_{r\to 1^{-}}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}}{\left(1-(1-2e^{-t})r\right)^{\alpha}}dt=\int_{0}^{\infty}e^{-\alpha t}dt=\frac{1}{\alpha}.$
Next, we derive the upper bound.
Let $f\in H_{\alpha}^{\infty}$ with $0<\alpha\leq\frac{1}{2}$. Using the estimate $|\phi_{t}^{\prime}(z)|=|\frac{e^{-t}}{(1-(1-e^{-t})z)^{2}}|\leq e^{t}$ and the Schwarz-Pick lemma, we obtain that
$\|S_{t}(f)\|_{H_{\alpha}^{\infty}}$ $=\sup_{z\in\mathbb{D}}|S_{t}(f)(z)|(1-|z|^{2})^{\alpha}$
$=\sup_{z\in\mathbb{D}}\sqrt{e^{-t}}|\phi_{t}^{\prime}(z)|^{1/2}|f(\phi_{t}(z))|(1-|z|^{2})^{\alpha}$
$=\sup_{z\in\mathbb{D}}\sqrt{e^{-t}}|\phi_{t}^{\prime}(z)|^{1/2-\alpha}|\phi_{t}^{\prime}(z)|^{\alpha}|f(\phi_{t}(z))|(1-|z|^{2})^{\alpha}$
$\leq\sqrt{e^{-t}}(e^{t})^{1/2-\alpha}\sup_{z\in\mathbb{D}}|f(\phi_{t}(z))|(1-|z|^{2})^{\alpha}|\phi_{t}^{\prime}(z)|^{\alpha}$
$\leq e^{-\alpha t}\sup_{z\in\mathbb{D}}|f(\phi_{t}(z))|(1-|\phi_{t}(z)|^{2})^{\alpha}$
$\leq e^{-\alpha t}\|f\|_{H_{\alpha}^{\infty}}.$
Then,
$\|\mathcal{C}(f)\|_{H_{\alpha}^{\infty}}$ $=\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\left|\int_{0}^{\infty}S_{t}(f)(z)dt\right|$
$\leq\int_{0}^{\infty}\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}|S_{t}(f)(z)|dt$
$\leq\int_{0}^{\infty}\|S_{t}(f)\|_{H_{\alpha}^{\infty}}dt$
$\leq\int_{0}^{\infty}e^{-\alpha t}\|f\|_{H_{\alpha}^{\infty}}dt$
$=\frac{1}{\alpha}\|f\|_{H_{\alpha}^{\infty}}.$
Therefore, for $0<\alpha\leq\frac{1}{2}$,
$\|\mathcal{C}\|_{H_{\alpha}^{\infty}\rightarrow H^{\infty}_{\alpha}}=\frac{1}{\alpha}.$
This completes the proof of the theorem.
4. Norm estimates of the Cesàro operator $\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow H^{\infty}_{\alpha}}$
Since $H^{\infty}_{\alpha,\log}\subsetneqq H^{\infty}_{\alpha}$, for $0<\alpha<1$, and considering that the norm of the Cesàro operator $\mathcal{C}$ on the Korenblum space is bounded in this range, we can conclude that $\mathcal{C}$ is a bound operator acting from $H^{\infty}_{\alpha,\log}$ into $H^{\infty}_{\alpha}$. In this section, we aim to derive norm estimates for the Cesàro operator as it acts from $H^{\infty}_{\alpha,\log}$ into $H^{\infty}_{\alpha}$ for $0<\alpha<1$.
Theorem 4.1.
For $0<\alpha<1$, then
$\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow H^{\infty}_{\alpha}}=\sup_{0\leq r<1}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{re^{-t}}{1-(1-e^{-t})r}\right)^{2}}}dt,$
and
$\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow H^{\infty}_{\alpha}}\geq\frac{1}{\frac{1}{\alpha}+\log 2}.$
Proof.
Let $0<\alpha<1$ and $z\in\mathbb{D}$. Define
$f_{\alpha}(z)=\frac{1}{(1-z^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-z^{2}}}.$
By a simple calculation, we see that $g(x)=x^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{x}$ is monotonically increasing in $(0,2)$. Since $0\leq|1-z^{2}|\leq 2$, we obtain that
$\|f_{\alpha}\|_{H^{\infty}_{\alpha,\log}}$ $=\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-|z|^{2}}\left|\frac{1}{(1-z^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-z^{2}}}\right|$
$\leq\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-|z|^{2}}\frac{1}{(1-|z|^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-|z|^{2}}}$
$=\sup_{0\leq r<1}(1-r^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-r^{2}}\frac{1}{(1-r^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-r^{2}}}$
$=1.$
Since
$\lim\limits_{r\rightarrow 1}|f_{\alpha}(z)|(1-r^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-r^{2}}=1,$
we conclude $\|f_{\alpha}\|_{H^{\infty}_{\alpha,\log}}=1$.
The weighted composition operator $S_{t}$ applied to a function $f_{\alpha}$ can be written as
$S_{t}f_{\alpha}(z)$ $=\frac{e^{-t}}{1-(1-e^{-t})z}f(\frac{e^{-t}z}{1-(1-e^{-t})z})$
$=\frac{e^{-t}}{1-(1-e^{-t})z}\frac{1}{(1-(\frac{e^{-t}z}{1-(1-e^{-t})z})^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-(\frac{e^{-t}z}{1-(1-e^{-t})z})^{2}}}$
$=\frac{e^{-t}\left(1-(1-e^{-t})z\right)^{2\alpha-1}}{(1-z)^{\alpha}\left(1-(1-2e^{-t})z\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{e^{-t}z}{1-(1-e^{-t})z}\right)^{2}}}$
Since $\mathcal{C}f_{\alpha}(z)=\int_{0}^{\infty}S_{t}f_{\alpha}(z)dt$ and $\|f_{\alpha}\|_{H^{\infty}_{\alpha,\log}}=1,$ we have
$\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow H^{\infty}_{\alpha}}$ $\geq\|\mathcal{C}f_{\alpha}\|_{H^{\infty}_{\alpha}}=\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}|\mathcal{C}f_{\alpha}(z)|$
$=\sup_{z\in\mathbb{D}}\int_{0}^{\infty}(1-|z|^{2})^{\alpha}\left|\frac{e^{-t}\left(1-(1-e^{-t})z\right)^{2\alpha-1}}{(1-z)^{\alpha}\left(1-(1-2e^{-t})z\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{e^{-t}z}{1-(1-e^{-t})z}\right)^{2}}}\right|dt$
$\geq\sup_{0\leq r<1}\int_{0}^{1}(1-r^{2})^{\alpha}\frac{e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}}{(1-r)^{\alpha}\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{re^{-t}}{1-(1-e^{-t})r}\right)^{2}}}dt$
$=\sup_{0\leq r<1}\int_{0}^{1}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{re^{-t}}{1-(1-e^{-t})r}\right)^{2}}}dt.$ (4.1)
On the other hand, we have that
$|S_{t}f(z)|=|w_{t}(z)f(\phi_{t}(z))|$
$=\lvert w_{t}(z)\rvert\frac{1}{(1-|\phi_{t}(z)|^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-|\phi_{t}(z)|^{2}}}(1-|\phi_{t}(z)|^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-|\phi_{t}(z)|^{2}}|f(\phi_{t}(z))|$
$\leq\lvert w_{t}(z)\rvert\frac{1}{(1-|\phi_{t}(z)|^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-|\phi_{t}(z)|^{2}}}\|f\|_{H^{\infty}_{\alpha,\log}},$ (4.2)
since $|w_{t}(z)|=\frac{e^{-t}}{|1-(1-e^{-t})z|}\leq\frac{e^{-t}}{1-(1-e^{-t})|z|}$, $|\phi_{t}(z)|=\frac{|e^{-t}z|}{|1-(1-e^{-t})z|}\leq\frac{e^{-t}|z|}{1-(1-e^{-t})|z|}$ and the monotonicity of $g(x)$, we obtain
$|S_{t}f(z)|\leq w_{t}(|z|)\frac{1}{(1-\phi_{t}(|z|)^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\phi_{t}(|z|)^{2}}}\|f\|_{H^{\infty}_{\alpha,\log}}.$
Therefore, from inequality (4) we deduce
$\|\mathcal{C}f\|_{H^{\infty}_{\alpha}}$ $=\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}|\int_{0}^{\infty}S_{t}f(z)dt|$
$\leq\sup_{0\leq r<1}(1-r^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-t}}{1-(1-e^{-t})r}\frac{1}{(1-(\frac{e^{-t}r}{1-(1-e^{-t})r})^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-(\frac{e^{-t}r}{1-(1-e^{-t})r})^{2}}}dt\|f\|_{H^{\infty}_{\alpha,\log}}$
$=\sup_{0\leq r<1}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{re^{-t}}{1-(1-e^{-t})r}\right)^{2}}}dt\|f\|_{H^{\infty}_{\alpha,\log}}.$ (4.3)
Hence,
$\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow H^{\infty}_{\alpha}}=\sup_{0\leq r<1}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{re^{-t}}{1-(1-e^{-t})r}\right)^{2}}}dt<\infty.$
Since
$\sup_{0\leq r<1}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{re^{-t}}{1-(1-e^{-t})r}\right)^{2}}}dt$
$\geq\sup_{r=0}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{re^{-t}}{1-(1-e^{-t})r}\right)^{2}}}dt=\frac{1}{\log 2e^{\frac{1}{\alpha}}}\int_{0}^{\infty}e^{-t}dt$
Thus,
$\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow H^{\infty}_{\alpha}}\geq\frac{1}{\log 2+\frac{1}{\alpha}}.$
This completes the proof. ∎
5. Norm estimates of the Cesàro operator $\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow H^{\infty}_{\alpha,\log}}$
In this section, we calculate the norm of the Cesàro operator $\mathcal{C}$ acting on the logarithmically weighted Korenblum space $H^{\infty}_{\alpha,\log}$.
Theorem 5.1.
For $0<\alpha<1$, then the Cesàro operator $\mathcal{C}$ is bounded on the logarithmically weighted Korenblum space $H^{\infty}_{\alpha,\log}$. Moreover, the norm of $\mathcal{C}$ satisfies the following equations:
$\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow H^{\infty}_{\alpha,\log}}=\sup_{0\leq r<1}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}\log\frac{2e^{\frac{1}{\alpha}}}{1-r^{2}}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{e^{-t}r}{1-(1-e^{-t})r}\right)^{2}}}dt.$
and
$\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow{H^{\infty}_{\alpha,\log}}}\geq\frac{1}{\alpha}.$
Proof.
Let
$f_{\alpha}(z)=\frac{1}{(1-z^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-z^{2}}}.$
From the proof of Theorem 4.1, we know that $\|f_{\alpha}\|_{H^{\infty}_{\alpha,\log}}=1$ and
$S_{t}f_{\alpha}(z)=\frac{e^{-t}\left(1-(1-e^{-t})z\right)^{2\alpha-1}}{(1-z)^{\alpha}\left(1-(1-2e^{-t})z\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{e^{-t}z}{1-(1-e^{-t})z}\right)^{2}}}.$
Since $\mathcal{C}f_{\alpha}(z)=\displaystyle\int_{0}^{\infty}S_{t}f_{\alpha}(z)dt$, we deduce
$\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow H^{\infty}_{\alpha,\log}}$ $\geq\|\mathcal{C}f_{\alpha}\|_{H^{\infty}_{\alpha,\log}}$
$=\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-|z|^{2}}|\mathcal{C}f_{\alpha}(z)|$
$\geq\sup_{0\leq r<1}(1-r^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-r^{2}}\mathcal{C}f_{\alpha}(r)$
$=\sup_{0\leq r<1}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}\log\frac{2e^{\frac{1}{\alpha}}}{1-r^{2}}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-(\frac{e^{-t}r}{1-(1-e^{-t})r})^{2}}}dt.$
On the other hand, using the estimate
$|S_{t}f(z)|\leq w_{t}(|z|)\frac{1}{(1-\phi_{t}(|z|)^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\phi_{t}(|z|)^{2}}}\|f\|_{H^{\infty}_{\alpha,\log}},$
we obtain that
$\|\mathcal{C}f\|_{H^{\infty}_{\alpha,\log}}$ $=\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-|z|^{2}}|\int_{0}^{\infty}S_{t}f(z)dt|$
$\leq\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-|z|^{2}}|\int_{0}^{\infty}w_{t}(|z|)\frac{1}{(1-\phi_{t}(|z|)^{2})^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\phi_{t}(|z|)^{2}}}dt\|f\|_{H^{\infty}_{\alpha,\log}}$
$\leq\sup_{0\leq r<1}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}\log\frac{2e^{\frac{1}{\alpha}}}{1-r^{2}}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-(\frac{e^{-t}r}{1-(1-e^{-t})r})^{2}}}dt\|f\|_{H^{\infty}_{\alpha,\log}}.$
Therefore,
$\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow H^{\infty}_{\alpha,\log}}=\sup_{0\leq r<1}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}\log\frac{2e^{\frac{1}{\alpha}}}{1-r^{2}}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{e^{-t}r}{1-(1-e^{-t})r}\right)^{2}}}dt.$
Now, observe that
$\lim_{r\rightarrow 1^{-}}\frac{\log\frac{2e^{\frac{1}{\alpha}}}{1-r^{2}}}{\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{e^{-t}r}{1-(1-e^{-t})r}\right)^{2}}}=1,$
Hence,
$\|\mathcal{C}\|_{H^{\infty}_{\alpha,\log}\rightarrow{H^{\infty}_{\alpha,\log}}}$ $\geq\lim_{r\rightarrow 1^{-}}\int_{0}^{\infty}\frac{(1+r)^{\alpha}e^{-t}\left(1-(1-e^{-t})r\right)^{2\alpha-1}\log\frac{2e^{\frac{1}{\alpha}}}{1-r^{2}}}{\left(1-(1-2e^{-t})r\right)^{\alpha}\log\frac{2e^{\frac{1}{\alpha}}}{1-\left(\frac{e^{-t}r}{1-(1-e^{-t})r}\right)^{2}}}dt$
$=\int_{0}^{\infty}e^{-\alpha t}dt$
$=\frac{1}{\alpha}$
[…]
From [23], we know that the Cesàro operator $\mathcal{C}$ is bounded on the $\alpha$-Bloch space $\mathcal{B}^{\alpha}$ if and only if $1<\alpha<\infty$. Here, we provide an upper bound and a lower bound for the norm of the Cesàro operator when $\alpha>1$.
Lemma 6.1.
[12, Lemma 5.1] Let $f\in\mathcal{B}^{\alpha}$, then
[…]
Theorem 6.1.
For the Cesàro operator $\mathcal{C}$ acting on the space $\mathcal{B}^{\alpha}$, the following upper bounds hold for its norm:
$\|\mathcal{C}\|_{\mathcal{B}^{\alpha}\to\mathcal{B}^{\alpha}}\leq\begin{cases}\max\left\{A,\dfrac{2^{\alpha}}{\alpha-1}\right\},&\text{for }1<\alpha\leq 2,\\[20.00003pt] \max\left\{A,2^{\alpha}\dfrac{2^{\alpha}-\alpha-1}{(\alpha-1)^{2}}\right\},&\text{for }\alpha>2,\end{cases}$
where
$A=1+\left(\frac{2}{2\alpha-1}\right)^{2\alpha-1}\alpha^{\alpha}(\alpha-1)^{\alpha-1}.$
Proof.
Consider the upper bound of $\|\mathcal{C}\|_{\mathcal{B}^{\alpha}\rightarrow\mathcal{B}^{\alpha}}$. Let $f\in\mathcal{B}^{\alpha}$, we have that
$(\mathcal{C}f)^{\prime}(z)$ $=\int_{0}^{1}\frac{tf^{\prime}(tz)}{1-tz}dt+\int_{0}^{1}\frac{tf(tz)}{(1-tz)^{2}}dt$
$=\frac{1}{z^{2}}\int_{0}^{z}\frac{\xi f^{\prime}(\xi)}{1-\xi}+\frac{\xi f(\xi)}{(1-\xi)^{2}}d\xi,\quad z\in\mathbb{D}.$
We can change the path of integration to
$\xi=\phi_{t}(z)=\frac{e^{-t}z}{1-(1-e^{-t})z},\quad 0\leq t<\infty.$
Therefore, we obtain that
$(\mathcal{C}f)^{\prime}(z)=\int_{0}^{\infty}\frac{e^{-2t}}{1-\left(1-e^{-t}\right)z}\left[\frac{f^{\prime}(\phi_{t}(z))}{1-\left(1-e^{-t}\right)z}+\frac{f(\phi_{t}(z))}{1-z}\right]dt$
Then
$\|\mathcal{C}f(z)\|_{\mathcal{B}^{\alpha}}$ $=|\mathcal{C}f(0)|+\sup_{z\in\mathbb{D}}|\mathcal{C}f^{\prime}(z)|(1-|z|^{2})^{\alpha}$
$\leq|f(0)|+\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-2t}}{|1-(1-e^{-t})z|}\left(\frac{|f^{\prime}(\phi_{t}(z))|}{|1-(1-e^{-t})z|}+\frac{|f(\phi_{t}(z))|}{|1-z|}\right)dt.$
Since $|1-z|\geq 1-|z|=1-r$, $|f^{\prime}(\varphi_{t}(z))|\leq\frac{\|f\|_{\alpha^{*}}}{\left(1-|\phi_{t}(z)|^{2}\right)^{\alpha}}$, we can apply Lemma 6.1 to obtain the following estimate,
$\|\mathcal{C}f(z)\|_{\mathcal{B}^{\alpha}}$ $\leq|f(0)|+\sup_{0\leq|z|<1}(1-|z|^{2})^{\alpha}\int_{0}^{\infty}[\frac{e^{-2t}\|f\|_{\alpha^{*}}}{|1-(1-e^{-t})z|^{2}(1-|\phi_{t}(z)|^{2})^{\alpha}}$
$+\frac{e^{-2t}\left[(1-|\phi_{t}(z)|)^{1-\alpha}-1\right]\|f\|_{\alpha^{*}}}{(\alpha-1)(1-|z|)|1-(1-e^{-t})z|}+\frac{e^{-2t}|f(0)|}{(1-|z|)|1-(1-e^{-t})z|}]dt$
$\leq\left(1+\sup_{0\leq r<1}(1-r^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-2t}}{(1-r)(1-(1-e^{-t})r)}dt\right)|f(0)|$
$+\sup_{0\leq|z|<1}(1-|z|^{2})^{\alpha}\int_{0}^{\infty}[\frac{e^{-2t}}{|1-(1-e^{-t})z|^{2}(1-|\phi_{t}(z)|^{2})^{\alpha}}$
$+\frac{e^{-2t}\left[(1-|\phi_{t}(z)|)^{1-\alpha}-1\right]}{(\alpha-1)(1-|z|)|1-(1-e^{-t})z|}]\|f\|_{\alpha^{*}}dt\overset{def}{=}|f(0)|\cdot I+\|f\|_{\alpha^{*}}\cdot II.$
We estimate that
$I=1+\sup_{0\leq r<1}(1-r^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-2t}}{(1-r)(1-(1-e^{-t})r)}dt.$
$\leq 1+\sup_{0\leq r<1}(1+r)^{\alpha}(1-r)^{\alpha-1}\ \int_{0}^{\infty}e^{-t}dt$
$=1+\sup_{0\leq r<1}(1+r)^{\alpha}(1-r)^{\alpha-1}.$
Define the function
$f(r)=(1+r)^{\alpha}(1-r)^{\alpha-1}.$
A straightforward calculation shows that $f(r)$ attains its maximum when $r=\frac{1}{2\alpha-1}$. Therefore,
$1+\sup_{0\leq r<1}(1-r^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-2t}}{(1-r)(1-(1-e^{-t})r)}dt\leq 1+\left(\frac{2}{2\alpha-1}\right)^{2\alpha-1}\alpha^{\alpha}(\alpha-1)^{\alpha-1}\overset{def}{=}A.$ (6.1)
To bound $II$, we obtain that
$II=\sup_{0\leq|z|<1}(1-|z|^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-2t}}{|1-(1-e^{-t})z|^{2}(1-|\phi_{t}(z)|^{2})^{\alpha}}+\frac{e^{-2t}\left[(1-|\phi_{t}(z)|)^{1-\alpha}-1\right]}{(\alpha-1)(1-|z|)|1-(1-e^{-t})z|}dt.$
$\leq\sup_{0\leq|z|<1}(1-|z|^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-2t}}{|1-(1-e^{-t})z|^{2}(1-|\phi_{t}(z)|)^{\alpha}}+\frac{e^{-2t}(1-|\phi_{t}(z)|)^{1-\alpha}}{(\alpha-1)(1-|z|)|1-(1-e^{-t})z|}dt$
$=\sup_{0\leq|z|<1}(1-|z|^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-2t}|1-(1-e^{-t})z|^{\alpha-2}}{\left(|1-(1-e^{-t})z|-e^{-t}|z|\right)^{\alpha}}+\frac{e^{-2t}|1-(1-e^{-t})z|^{\alpha-2}}{(\alpha-1)(1-|z|)\left(|1-(1-e^{-t})z|-e^{-t}|z|\right)^{\alpha-1}}dt$
$\leq\sup_{0\leq|z|<1}(1-|z|^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-2t}|1-(1-e^{-t})z|^{\alpha-2}}{(1-|z|)^{\alpha}}+\frac{e^{-2t}|1-(1-e^{-t})z|^{\alpha-2}}{(\alpha-1)(1-|z|)^{\alpha}}dt$
$=\sup_{0\leq|z|<1}(1+|z|)^{\alpha}\frac{\alpha}{\alpha-1}\int_{0}^{\infty}e^{-2t}|1-(1-e^{-t})z|^{\alpha-2}dt.$
Now, we consider the case $1<\alpha\leq 2$. We estimate the supremum as follows:
$\sup_{0\leq|z|<1}(1+|z|)^{\alpha}\frac{\alpha}{\alpha-1}\int_{0}^{\infty}e^{-2t}|1-(1-e^{-t})z|^{\alpha-2}dt$
$\leq\sup_{0\leq r<1}(1+r)^{\alpha}\frac{\alpha}{\alpha-1}\int_{0}^{\infty}\frac{e^{-2t}}{(1-(1-e^{-t})r)^{2-\alpha}}dt$
$=2^{\alpha}\frac{\alpha}{\alpha-1}\int_{0}^{\infty}\frac{e^{-2t}}{e^{-(2-\alpha)t}}dt$
$=\frac{2^{\alpha}}{\alpha-1}$ (6.2)
Next, we consider the case $\alpha>2$. A similar estimation yields:
$\sup_{0\leq|z|<1}(1+|z|)^{\alpha}\frac{\alpha}{\alpha-1}\int_{0}^{\infty}e^{-2t}|1-(1-e^{-t})z|^{\alpha-2}dt$
$\leq\sup_{0\leq r<1}(1+r)^{\alpha}\frac{\alpha}{\alpha-1}\int_{0}^{\infty}\frac{e^{-2t}}{(1-(e^{-t}-1)r)^{2-\alpha}}dt$
$=2^{\alpha}\frac{\alpha}{\alpha-1}\int_{0}^{\infty}\frac{e^{-2t}}{(2-e^{-t})^{2-\alpha}}dt$
$=2^{\alpha}\frac{2^{\alpha}-\alpha-1}{(\alpha-1)^{2}}$ (6.3)
Combining (6.1), (6) and (6), we finally complete the proof.
Theorem 6.2.
For $1<\alpha<\infty$, we have
$\|\mathcal{C}\|_{\mathcal{B}^{\alpha}\rightarrow\mathcal{B}^{\alpha}}\geq\frac{3}{2}$
Proof.
Let $\alpha\not=1$ and $z\in\mathbb{D}$.Define
$f_{\alpha}=1$
we have the estimate
$\|f_{\alpha}\|_{\mathcal{B}^{\alpha}}=1$
According to (2.3), we obtain that
$\|\mathcal{C}\|_{\mathcal{B}^{\alpha}\rightarrow\mathcal{B}^{\alpha}}$ $\geq\frac{\|\mathcal{C}f_{\alpha}\|_{\mathcal{B}^{\alpha}}}{\|f_{\alpha}\|_{\mathcal{B}^{\alpha}}}$
$=|\mathcal{C}f(0)|+\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}|\mathcal{C}f_{\alpha}^{\prime}(z)|$
$=1+\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\big|\int_{0}^{\infty}\frac{e^{-t}(1-e^{-t})}{(1-(1-e^{-t})z)^{2}}dt\big|$
$\geq 1+\sup_{0\leq r<1}(1-r^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-t}(1-e^{-t})}{(1-(1-e^{-t})r)^{2}}dt$
$=1+\lim\limits_{r\rightarrow 0}(1-r^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-t}(1-e^{-t})}{(1-(1-e^{-t})r)^{2}}dt$
$=1+\int_{0}^{\infty}e^{-t}(1-e^{-t})dt$
$=1+\frac{1}{2}$
$=\frac{3}{2}.$
At this point, we have completed all the proof. ∎
7. Norm of Cesàro operator $\|\mathcal{C}\|_{H^{\infty}\rightarrow\mathcal{B}^{\alpha}}$
In this section we offer both the lower and upper bounds of the norm of the Cesàro operator from the Hardy space $H^{\infty}$ to $\alpha$-Bloch spaces $\mathcal{B}^{\alpha}$ with $\alpha\geq 1$.
Theorem 7.1.
For $\alpha\geq 1$, we obtain that
$3$ $\leq\|\mathcal{C}\|_{H^{\infty}\rightarrow\mathcal{B}^{\alpha}}\leq 4~~~\mbox{if}~~~\alpha=1,$
$\frac{3}{2}$ $\leq\|\mathcal{C}\|_{H^{\infty}\rightarrow\mathcal{B}^{\alpha}}\leq 1+3\cdot 2^{\alpha-1}~~~\mbox{if}~~~\alpha>1.$
Moreover, $\mathcal{C}$ is not bounded from $H^{\infty}$ to $\mathcal{B}^{\alpha}$ when $0<\alpha<1$.
Proof.
let $f=1\in H^{\infty}$, then
$\|f\|_{\infty}=1,~~\mathcal{C}(1)(z)=\frac{1}{z}\log\frac{1}{1-z},~~\mathcal{C}(1)(0)=1.$
We have that
$\|\mathcal{C}\|_{H^{\infty}\rightarrow\mathcal{B}^{\alpha}}$ $\geq|\mathcal{C}(f)(0)|+\|\mathcal{C}(f)(z)\|_{*}$
$=1+\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}|\mathcal{C}(1)^{\prime}(z)|$
$=1+\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\big|\frac{1}{z(1-z)}-\frac{1}{z^{2}}\log\frac{1}{1-z}\big|$
$\geq 1+\sup_{0\leq r<1}(1-r^{2})^{\alpha}\big|\frac{1}{r(1-r)}-\frac{1}{r^{2}}\log\frac{1}{1-r}\big|.$
Let $f(r)=\displaystyle(1-r^{2})^{\alpha}\big|\frac{1}{r(1-r)}-\frac{1}{r^{2}}\log\frac{1}{1-r}\big|$. By a simple calculation, we have that
$\sup_{0\leq r<1}f(r)$ $\geq\lim_{r\to 0^{+}}f(r)=+\infty~~~\mbox{if}~~~0<\alpha<1,$
$\sup_{0\leq r<1}f(r)$ $=\lim_{r\to 1}f(r)=2~~~\mbox{if}~~~\alpha=1.$
Therefore, $\mathcal{C}$ is not bounded from $H^{\infty}$ to $\mathcal{B}^{\alpha}$ when $0<\alpha<1$. For the case $\alpha>1$, similar to the proof of Theorem 6.2, we have that
$\|\mathcal{C}\|_{H^{\infty}\rightarrow\mathcal{B}^{\alpha}}$ $\geq|\mathcal{C}(1)(0)|+\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}|\mathcal{C}(1)^{\prime}(z)|$
$=1+\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}\big|\int_{0}^{\infty}\frac{e^{-t}(1-e^{-t})}{(1-(1-e^{-t})z)^{2}}dt\big|$
$\geq 1+\sup_{0\leq r<1}(1-r^{2})^{\alpha}\int_{0}^{\infty}\frac{e^{-t}(1-e^{-t})}{(1-(1-e^{-t})r)^{2}}dt$
$=\frac{3}{2}.$
This completes the proof of the lower bound. It remains to prove the upper bound for the case $\alpha\geq 1$.
Let $f\in H^{\infty}$. Then
${}\mathcal{C}(f)^{\prime}(z)=\int_{0}^{1}\frac{tf^{\prime}(tz)}{1-tz}dt+\int_{0}^{1}\frac{tf(tz)}{(1-tz)^{2}}dt,\quad z\in\mathbb{D}.$ (7.1)
By [24, Proposition 5.1], we have that
${}(1-|z|^{2})|f^{\prime}(z)|\leq\|f\|_{\infty}.$ (7.2)
for all $f\in H^{\infty}$ and $z\in\mathbb{D}$.
Using (7.1) and (7.2), we obtain that
$\|\mathcal{C}f\|_{\mathcal{B}^{\alpha}}$ $=|\mathcal{C}f(0)|+\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}|(\mathcal{C}f)^{\prime}(z)|$
$=|\int_{0}^{1}f(0)dt|+\sup_{z\in\mathbb{D}}(1-|z|^{2})^{\alpha}|(\mathcal{C}f)^{\prime}(z)|$
$=\left|\int_{0}^{1}f(0)dt\right|+\sup_{z\in\mathbb{D}}\left(1-|z|^{2}\right)^{\alpha}\left|\int_{0}^{1}\frac{tf^{\prime}(tz)}{1-tz}dt+\int_{0}^{1}\frac{tf(tz)}{(1-tz)^{2}}dt\right|$
$\leq\left(1+\sup_{0\leq r<1}\left(1-r^{2}\right)^{\alpha}\int_{0}^{1}\frac{2t+rt^{2}}{(1-tr)^{2}(1+tr)}dt\right)\|f\|_{\infty}$
$=\left(1+\sup_{0<r<1}\frac{(1-r^{2})^{\alpha}}{r^{2}}\left\{\frac{3r}{2(1-r)}-\frac{1}{4}\ln\frac{1+r}{(1-r)^{5}}\right\}\right)\|f\|_{\infty}.$
Let $h(r)=\displaystyle\frac{(1-r^{2})}{r^{2}}\left\{\frac{3r}{2(1-r)}-\frac{1}{4}\ln\frac{1+r}{(1-r)^{5}}\right\},0<r<1.$ We find that it can be shown that
$h(r)=1+r+\sum_{n=2}^{\infty}\left(\frac{5}{2n(n+2)}+\frac{(-1)^{n-1}}{2n(n+2)}\right)r^{n},$
then
$h^{\prime}(r)=1+\sum_{n=2}^{\infty}\left(\frac{5}{2(n+2)}+\frac{(-1)^{n-1}}{2(n+2)}\right)r^{n-1}>0.$
It implies that $h(r)$ is increasing and $\sup_{0<r<1}h(r)=\lim_{r\rightarrow 1^{-}}h(r)=3$.
Hence, we obtain that
$\|\mathcal{C}f\|_{\mathcal{B}^{\alpha}}\leq 1+\sup_{0<r<1}h(r)(1-r^{2})^{\alpha-1}\leq 1+3\cdot 2^{\alpha-1}.$
We finish the proof of the theorem.